Grade 11Discuss with colleagues and IITians

in any triangle ABC prove that 1/sin^nA+1/sin^nB+1/sin^nC>=3*2^n

S.Shrihari the genius
14 Years agoGrade 11

10 Answers

Swapnil Saxena
14 Years ago

The question is incorrect. Try it for n=1 and A= B=C= 60

Then the equation would be 1/sinn + 1/sinnB + 1/sinnC = 3 * (2/root(3))^1 =  root(3) 2 which is no doubt smaller than 3* 2^n

This equation is valid only for n < 0

Proof:

Using AM- GM inequalitites 

(1/sinn + 1/sinnB + 1/sinnC)/3 >= (1/sinn * 1/sinnB * 1/sinnC )1/3

With the equality existing only when 1/sinn = 1/sinnB = 1/sinnC

This can only be possible when sin A = sin B = sin C = 60

Then the minimum value of 1/sinn + 1/sinnB + 1/sinnC is ((2/root(3))3n )1/3

= ((2/root(3))n

= ((2n/root(3)n)*3

= (2n/3n-2/2)

now if (n-2)/2 is less than -1 or we can say n < 0 then the question is no doubt correct.

The correct question would be 1/sinn + 1/sinnB + 1/sinnC>= (2n/3n-2/2) or

1/sinn + 1/sinnB + 1/sinnC>= (3*2n) where n<0


S.Shrihari the genius
14 Years ago

no....the question is for n>=1 and n belogs to N

Swapnil Saxena
14 Years ago

Try it yourself with A=B=C=60 and n=1,2,3,4,...

You would surely find that the question is errorenous. Just try ones and u wil find that the question is true only for n<0.  

S.Shrihari the genius
14 Years ago

but you have computed the answer taking only the equality into consideration

S.Shrihari the genius
14 Years ago

and it is not only for A=B=C=60.............should work for any triangle

Swapnil Saxena
14 Years ago

OK... Then try it for a combination of 90,60,30. Is that Ok...

The result is 3 + 2/root(3) which is surely smaller than 6(3*2^1)

Now iv proved it fr inequality. 

S.Shrihari the genius
14 Years ago

proof for inequality.....not a verification

Swapnil Saxena
14 Years ago

Shrihari,

Iv already given u a proof using Am-Gm inequalitites in the first reply but if thr is something which u cant understand u can ask me. Certainly i m able to clear ur confusions but fr tht u hv to clr me what is ur opinion fr ur question . wht do you think the mistake is. Just try to look at my first reply .

Swapnil Saxena
14 Years ago

Fr understading my reply, u should first read about Am-Gm inequalities. I think u will certianly get tht my proof not only cover the case with equal angle but also with unequal angles. Till now im inable to understand your prob...

S.Shrihari the genius
14 Years ago

Sir,

I have read about AM-GM and i understand your proof........thank you.......sir,